Determine the intervals on which the graph of the following curve is concave up/down: x=cos(t) , y=sin(2t) , on [0,2pi]

Please show work clearly, thanks!
Answer #1

accolt), = siult) (helo] no up dj = -sincti, dy =2cos (at) at dy dy/at an = Los(at) - drfat -Sint) -2 cos (24) Sinit) :dy dandy due 10 For critical points -2 tost) (+ )) (+) -24)=9 1+ 2SiY=0 ㅋ 420 S글 키크포,3플 Sint) t = 포 lo 욘 Jr fC스 에 ( )2 +) f(-) -w